OpenStudy (anonymous):

Xavier is riding on a Ferris wheel at the local fair. His height can be modeled by the equation H(t) = 20cosine of the quantity 1 over 15 times t + 30, where H represents the height of the person above the ground in feet at t seconds. Part 1: How far above the ground is Xavier before the ride begins? Part 2: How long does the Ferris wheel take to make one complete revolution? Part 3: Assuming Xavier begins the ride at the top, how far from the ground is the edge of the Ferris wheel, when Xavier's height above the ground reaches a minimum? You must show all work.

OpenStudy (anonymous):

I'll fan and Medal!!!

OpenStudy (anonymous):

@ganeshie8

OpenStudy (anonymous):

@dan815

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

@MarrBear

OpenStudy (amistre64):

any ideas?

OpenStudy (anonymous):

Well, for the first one, I think I need to substitute 0 in for t.

OpenStudy (anonymous):

I am rather confused on the units they want it in- am I supposed to use my calculator and find decimals? or do they want it in the equation, unsimplified?

OpenStudy (amistre64):

hmm, well lets draw a picture 

OpenStudy (amistre64):

since the amplitude is given as 20, the radius of the wheel is 20 its midline is given as +30, so its center is 30 feet off the ground. when we enter into the wheel, we enter in at the bottom ....

OpenStudy (amistre64):

but i may be reading the wording wrong since it is not in a mathmatical notation

OpenStudy (amistre64):

otherwise, t=0 sounds like a fine starting point :)

OpenStudy (anonymous):

One moment- I'll try to put it in mathmatics

OpenStudy (anonymous):

\[20 \cos \left( \frac{ \pi }{ 15 }t \right)+30\]

OpenStudy (anonymous):

Sorry- I'm very slow at that

OpenStudy (amistre64):

so theres a pi involved?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Does that mean they want radians?

OpenStudy (amistre64):

that was missing from the original if t=0, and t assumes a proper time frame. then it looks like letting t=0 is a good bet for part a

OpenStudy (amistre64):

and you are sure the +30 is on the outside of cosine?

OpenStudy (anonymous):

I copy and pasted the original- the formatting must have been messed up

OpenStudy (anonymous):

I'm positive- the problem shows the problem in mathematical notation

OpenStudy (amistre64):

a screen shot or picture to attach would really help with some of the confusion :)

OpenStudy (amistre64):

k, then if we let t=0, the we start out at 20+30, since cos(0)=1

OpenStudy (anonymous):

let me see...

OpenStudy (anonymous):

how do I screenshot again? I'm sorry

OpenStudy (amistre64):

you on windows or a mac?

OpenStudy (anonymous):

A chromebook

OpenStudy (amistre64):

on windows, there is a set of button that never get used. prnt scrn, scroll lock, and pause/break the prnt scrn button takes a picture of whatever is on your screen and can be pasted into paint

OpenStudy (amistre64):

a chrome book, havent a clue